The subject of this case study is “Jack”, a 52‐year‐old Caucasian male, former‐smoker (5 years ago), living and working in Hong Kong.    Jack has recently set up a consultancy company, having worked for many years within large organizations. This change is allowing him more time to focus on other aspects of his life (family and friends, health and wellbeing, study and travel) and to create a better work‐life balance.  Jack is running his business from home.   In the last four years Jack has become a long distance runner, competing in the London marathon in 2009 and the Paris marathon in 2010.  Without a marathon in his schedule, Jack is finding it hard to stay motivated to train.  His current height is 6.0, weight is 81kg, which is 4kg heavier than his racing weight of 77kg.  He confesses to not having changed his eating patterns despite not training or working with the same intensity as before.    Jack’s objective is to reduce his body mass by 5% and return to 77kg quickly but safely.  Jack certainly likes a beer or two in the evenings with his dinner, and he allergic to any NUT products and dairy.

Calculate the:

  1. BMR
  2. Calories per day before goals
  3. Calories per day after goals
  4. Protein Intake in Calories and Grams
  5. Fat intake in Calories and Grams
  6. Carbohydrate intake in Calories and Grams
  7. Look at what Jack has eaten yesterday. How far off is he with his goals in each category? (Carbs, Protein, Fat and Calories?
Answer #1

BMR can be calculated with miffin st jeor formula

ie, For men,

BMR=10W+6.25H-5A+5

calories / day before goals =(10*81)+(6.25*180)-(5*52)+5

=1680 calories/day

calories / day after goals=(10*77)+(6.25*180)-(5*52)+5

=1640 calories/day

For attaining a weightloss with 0.5kg/week,he should take

protien-110 grams/day (110*4=440cal/day)

carbohydrates-241 grams/day (241*4=964cal/day)

fat-51 grams/day( 51*9=459cal/day)

Jack could reach the goal of 77kg with this diet and exercises 3 times a week in 8 weeks